OpenStudy (anonymous):

find limit as x approaches 0 of (e^x + 5x)^(2/x)

OpenStudy (anonymous):

so i got the indeterminate form of 1^infinity what should ido next?

zepdrix (zepdrix):

\[\Large\rm \lim_{x\to0}(e^x+5x)^{2/x}\]Okkayyyyy. I guess we should logs.\[\LARGE\rm e^{\left[\ln\lim_{x\to0}(e^x+5x)^{2/x}\right]}\]

zepdrix (zepdrix):

Does that step make sense or not so much? If it's easier, we can write a "left side" as y, and then take the log of both sides.

zepdrix (zepdrix):

This is a good trick to get comfortable with though

OpenStudy (anonymous):

yes please continue :D

zepdrix (zepdrix):

Ignoring the base e for a moment, let's see what's happening up in the exponent.\[\Large\rm \ln\lim_{x\to0}(e^x+5x)^{2/x}\]Pass the log into the limit,\[\Large\rm \lim_{x\to0}\ln\left[(e^x+5x)^{2/x}\right]\]Applying rules of logs allows us to bring the 2/x outside,\[\Large\rm \lim_{x\to0}\frac{2}{x}\ln(e^x+5x)\]

zepdrix (zepdrix):

Let's write the x in the denominator, it might give us one of the indeterminate forms that we can work with,\[\Large\rm 2\lim_{x\to0}\frac{\ln(e^x+5x)}{x}\]

zepdrix (zepdrix):

So what is our stuff approaching now? :)

OpenStudy (anonymous):

u mean to the left or to right?

zepdrix (zepdrix):

Err I guess what I mean is.... Just plug in x=0 and see what indeterminate form we're getting. We should still be getting an indeterminate form.

OpenStudy (anonymous):

but x in the bottom is 0

zepdrix (zepdrix):

So the bottom is approaching 0. What is the top approaching?

OpenStudy (anonymous):

0 too?

OpenStudy (anonymous):

ln(1) = 0

zepdrix (zepdrix):

Ok good! So we're getting the indeterminate form 0/0. Which is one that allows us to use our old friend \(\Large\rm L'Hospital!\)

OpenStudy (anonymous):

derivative right

zepdrix (zepdrix):

Yes derivative of top and bottom, separately.

OpenStudy (anonymous):

for top I got (e^x +5) / (5x + e^x) for bottom I got 1/x^2

OpenStudy (anonymous):

is this right

zepdrix (zepdrix):

Woops let's fix the bottom a sec. The bottom is x. `It is not` 1/x.

OpenStudy (anonymous):

yeah i keep make mistake over and over

zepdrix (zepdrix):

>.< heh

OpenStudy (anonymous):

so its 1

zepdrix (zepdrix):

Ok great, so we'll simplify the fraction down, \[\Large\rm 2\lim_{x\to0}\frac{e^x+5}{e^x+5x}\]

zepdrix (zepdrix):

So what are we getting now? Still indeterminate form or no?

OpenStudy (anonymous):

12

zepdrix (zepdrix):

Good! And remember that we had that base of e right??

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

e^12

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

wohoho thank you so much

zepdrix (zepdrix):

np :3